The denominator of a fraction is one more than the numerator. If the numerator is increased by three, the resulting fraction is one more than the original fractional. Find the original fraction.

Answer by rothauserc(4718)

let x be the numerator of our fraction, then
x / (x+1) = ((x+3) / (x+1)) +1
multiply both sides of = by (x+1)
x = (x+3) + (x+1)
combine like terms
x = 2x +4
x = -4
check our answer in original equation
-4 / (-4+1) = ((-4+3) / (-4+1)) +1
combine like terms
4/3 = 1/3 + 3/3 = 4/3
our answer checks and the original fraction is 4/3